If cos θ=-5/13 and 0≤x ≤2 π, determine:
a) tan θ
b) sec θ
c) the possible values of θ to the nearest tenth

Answer by stanbon(48558)

If cos θ=-5/13 and 0≤x ≤2 π, determine:
Since cos = x/5, x = -5 and r = 13
Then y = sqrt(r^2-x^2) = 12 or -12
----------------
a) tan θ = y/x = +12/13 or -12/13
-------------------
b) sec θ = r/x = 13/-5
---------------------
c) the possible values of θ to the nearest tenth
theta = inversetan(-12/13) = 180 - 42.7 = 137.3
or
theta = inverseta(12/13) = 122.7
=====================================
Cheers,
Stan H.